# VoxCity: A Seamless Framework for Open Geospatial Data Integration, Grid-Based Semantic 3D City Model Generation, and Urban Environment Simulation

### v1

@@ 4.3.2 OBJ for 3D modeling and rendering
The 3D array from VoxCity is converted into surface polygons of the voxel model and then formatted according to the OBJ structure. Colors corresponding to voxel classes are assigned to polygons, and the color palettes are saved as MTL files. This allows 3D modeling software to automatically apply the assigned surface colors when loading OBJ files and helps preserve semantic information that is not native to OBJ [161].
@@ 4.4 Visualization
VoxCity provides a built-in function for 3D visualization. Figure 11 shows example outputs produced by this functionality. Users can visualize generated 3D city models as well as simulation results. For instance, Figure 11a, b, c, and d indicate the yearly cumulative global irradiance on building surfaces, the ground-level sky view index, the ground-level green view index, and the ground-level landmark visibility, respectively. Additionally, users can specify projection types from two options: ‘orthographic’ and ‘perspective’, adjust the zoom factor, and change the camera position and angle.
@@
5 Limitations and future direction
@@ 5.1 Support and integration of existing semantic 3D city models
Second, in the current implementation, VoxCity provides four export formats (i.e., NPY, VOX, INX, and OBJ) to interface with downstream applications. Fostering data standardization and interoperability, in the future, we plan to implement additional export options (e.g., CityGML or CityJSON). For example, exporting 3D city models in standardized formats is of value for use cases that involve data exchange and sharing with practitioners or systems, which require compliance with established schemas. Offering CityGML or CityJSON as export options, VoxCity can enhance compatibility and ensure broader usability across various domains. Further, VoxCity incorporates semantic information during the generation process, preserving these attributes through standardized formats is important. Formats such as CityJSON are ideal for this purpose, providing a compact, structured, and easy-to-use approach to storing and managing semantic data [162]. Therefore, such properties can be retained and effectively communicated for downstream practices. For example, a recent study by Lei et al. [165] extended the CityJSON schema to accommodate human perception of architectural appearance in 3D buildings, enabling potential use cases. In this work, VoxCity features its novelty in automatic generation and tangible usability. Enabling standardized export can strengthen the role of VoxCity as a versatile tool, supporting the consistent interoperable development of 3D city models. However, challenges should be considered, such as the complexity of translating a voxel-based representation into a hierarchical structure, as well as computational and storage costs compared to current formats (e.g., NPY or OBJ).
@@ 5.2 Integration of building surface material information
Integrating building surface attributes, such as surface materials and window-to-wall ratio (WWR), to VoxCity is another promising direction for the next iteration of the package. VoxCity currently does not support the integration of building surface attributes, primarily due to limitations in data sources. To the best of our knowledge, no global open datasets currently provide surface attributes for individual buildings. While some buildings in OSM include material information, the proportion is relatively small [108]. However, building materials and WWR can significantly influence the surrounding microclimate and building energy consumption in various ways, including heat transfer, air leakage or ventilation, and offsetting daylighting demand [166, 167, 168]. This highlights the importance of incorporating material and WWR information in 3D city model generation for detailed urban environment simulations. A potential approach to address this gap is to use computer vision techniques on street view imagery to infer facade features.
@@
In light of this, we conducted a trial to incorporate (1) building surface materials and (2) WWR information into VoxCity’s 3D city model generation. Figure 12 provides an overview of the process. First, street view images were obtained from Mapillary for a target area in Houston, United States. OpenFACADES pipeline [169] is then employed to geolocate and detect buildings within the street view images, assigning them to their corresponding building footprints. Building on previous facade classification studies [170, 171, 172], we applied fine-tuned models to categorize the building materials depicted in these images [173, 174], with each footprint linked to multiple images. For WWR, window pixels were identified using the Grounded Segment Anything [175], and the percentage of window pixels within the building walls was calculated. Based on the inferred materials and WWR, we assigned categorical values to building voxels (e.g., 10 for glass, 11 for concrete, 12 for brick, and 13 for wood).
@@
5.3 Coverage of open geospatial data
In Section 4, we have demonstrated VoxCity’s capability to generate 3D models of a variety of cities worldwide. However, it cannot be used to generate models for every urban area, mostly because existing open datasets for building footprints and heights remain incomplete, and VoxCity does not incorporate all available data sources. Consequently, users may find it challenging to produce accurate models for regions beyond its current scope. We are addressing this challenge by continuously updating VoxCity and expanding the range of datasets it supports. Meanwhile, to broaden the utilization of building data, efforts to combine individual local datasets into comprehensive global collections—such as EUBUCCO [105]—should be further encouraged.
5.4 Incorporation of tree inventory
@@ 6 Conclusions
    By leveraging data from OSM, which is continuously updated, VoxCity helps its 3D city models remain current. Many open 3D city models are not regularly updated and can quickly become obsolete; VoxCity addresses this limitation by providing a more up-to-date representation of the urban form.
@@ Acknowledgements
We express our gratitude to the members of the NUS Urban Analytics Lab for the valuable discussions. This research has been supported by Takenaka Corporation. This research is part of the projects (i) Large-scale 3D Geospatial Data for Urban Analytics, which is supported by the National University of Singapore under the Start Up Grant; (ii) Multi-scale Digital Twins for the Urban Environment: From Heartbeats to Cities, which is supported by the Singapore Ministry of Education Academic Research Fund Tier 1; (iii) NUS Resilience and Growth Postdoctoral Fellowship – Smart Cities and Urban Analytics, which is funded by the National Research Foundation (NRF), Singapore; (iv) Development of a Multiscale Urban Microclimate Model for NUS Campus Thermal Environment (funded by NUS). We would like to thank the Singapore International Graduate Award (SINGA) scholarship provided by the Agency for Science, Technology, and Research (A*STAR), and NUS Research Scholarship and the President’s Graduate Fellowship (PGF) provided by NUS.
@@
The scripts developed for our method are openly available at a GitHub repository (https://github.com/kunifujiwara/VoxCity).

### v3

@@ 4.3.2 OBJ for 3D modeling and rendering
The 3D array from VoxCity is converted into surface polygons of the voxel model and then formatted according to the OBJ structure. Colors corresponding to voxel classes are assigned to polygons, and the color palettes are saved as Material Template Library (MTL) files. This allows 3D modeling software to automatically apply the assigned surface colors when loading OBJ files and helps preserve semantic information that is not native to OBJ [157]. Simulation results from VoxCity’s internal functions can also be exported as OBJ and MTL files.
@@ 4.4 Visualization
VoxCity provides a built-in function for 3D visualization. Figure 11 shows example outputs produced by this functionality. Users can visualize generated 3D city models as well as simulation results. For instance, Figure 11a, b, c, and d indicate the yearly cumulative global irradiance on building surfaces, the ground-level sky view index, the ground-level green view index, and the ground-level landmark visibility, respectively. Users can specify projection types from two options: ‘orthographic’ and ‘perspective’, adjust the zoom factor, and change the camera position and angle.
@@
5 Limitations and future directions
@@ 5.1 Support and integration of existing semantic 3D city models
Second, in the current implementation, VoxCity provides four export formats (i.e., PKL, VOX, INX, and OBJ) to interface with downstream applications. Fostering data standardization and interoperability, in the future, we plan to implement additional export options (e.g., CityGML or CityJSON). For example, exporting 3D city models in standardized formats is of value for use cases that involve data exchange and sharing with practitioners or systems, which require compliance with established schemas. By offering CityGML or CityJSON as export options, VoxCity can enhance compatibility and ensure broader usability across various domains. Further, VoxCity incorporates semantic information during the generation process; therefore, preserving these attributes through standardized formats is important. Formats such as CityJSON are ideal for this purpose, providing a compact, structured, and easy-to-use approach to storing and managing semantic data [158]. Therefore, such properties can be retained and effectively communicated for downstream practices. For example, a recent study by Lei et al. [161] extended the CityJSON schema to accommodate human perception of architectural appearance in 3D buildings, enabling potential use cases. In this work, VoxCity features its novelty in automatic generation and tangible usability. Enabling standardized export can strengthen the role of VoxCity as a versatile tool, supporting the consistent and interoperable development of 3D city models. However, challenges should be considered, such as the complexity of translating a voxel-based representation into a hierarchical structure, as well as computational and storage costs compared to current formats (e.g., PKL or OBJ).
@@ 5.2 Integration of building surface material information
Integrating building surface attributes, such as surface materials and window-to-wall ratio (WWR), into VoxCity is another promising direction for the next iteration of the package. VoxCity currently does not support the integration of building surface attributes, primarily due to limitations in data sources. To the best of our knowledge, no global open datasets currently provide surface attributes for individual buildings. While some buildings in OSM include material information, the proportion is relatively small [107]. However, building materials and WWR can significantly influence the surrounding microclimate and building energy consumption in various ways, including heat transfer, air leakage or ventilation, and offsetting daylighting demand [162, 163, 164]. This highlights the importance of incorporating material and WWR information in 3D city model generation for detailed urban environment simulations. A potential approach to address this gap is to use computer vision techniques on street view imagery to infer facade features.
@@
In light of this, we conducted a trial to incorporate (1) building surface materials and (2) WWR information into VoxCity’s 3D city model generation. Figure 12 provides an overview of the process. First, street view images were obtained from Mapillary for a target area in Houston, United States. OpenFACADES pipeline [165] was then employed to geolocate and detect buildings within the street view images, assigning them to their corresponding building footprints. Building on previous facade classification studies [166, 167, 168], we applied fine-tuned models to categorize the building materials depicted in these images [169, 170], with each footprint linked to multiple images. For WWR, window pixels were identified using the Grounded Segment Anything [171], and the percentage of window pixels within the building walls was calculated. Based on the inferred materials and WWR, we assigned categorical values to building voxels (e.g., 10 for glass, 11 for concrete, 12 for brick, and 13 for wood).
@@
5.3 Data availability and functionality constraints
In Section 4, we have demonstrated VoxCity’s capability to generate 3D models of a variety of cities worldwide. However, we acknowledge that data availability remains a critical challenge, and VoxCity does not incorporate all available data sources. Consequently, users may find it challenging to produce accurate models for regions beyond its current scope. Notwithstanding, comprehensive tools such as VoxCity play a crucial role in highlighting where data gaps remain, thereby supporting the development of targeted use cases. We are further addressing this challenge by continuously updating VoxCity and expanding the range of datasets it supports. Meanwhile, to broaden the utilization of building data, efforts to combine individual local datasets into comprehensive global collections—such as EUBUCCO [104]—should be further encouraged. Additionally, we have provided only qualitative comparisons of data sources in Section 3. To help users select the best data source options for their target areas, future work will conduct comparative experiments that provide quantitative comparisons of data sources in coverage and completeness.
@@ 5.3 Data availability and functionality constraints
VoxCity currently does not support the ability to edit data source files. For example, users cannot edit the arrangement of buildings, trees, and land cover from the original datasets. Therefore, users cannot conduct simulations to compare different design scenarios. To further enhance usability in urban studies and urban and architectural planning practices, future work will address this limitation by developing, for instance, a user interface to interactively edit data source files and voxel city models. Additionally, the current version does not support multi-dimensional data overlays, or integration with GIS platforms. These limitations represent important areas for future development of VoxCity.
5.4 Validity and reliability of generated models
The validity of the 3D city models generated by VoxCity depends on the accuracy of the input datasets. Each incorporated dataset contains its own uncertainty levels, as shown in Tables 1, 2, 4, and 5. For example, OB2.5DT (building height) has an MAE of approximately 1.5 m [26], META (tree canopy height) exhibits an MAE of 2.8 m [23], and FABDEM (terrain elevation) shows MAEs of 1–3 m depending on land cover [25]. Consequently, the accuracy of the generated 3D city models is constrained by these inherent uncertainties in the source data. Users are therefore encouraged to interpret simulation outputs—such as solar irradiance or view indices—within the expected uncertainty range of the input data. Nonetheless, VoxCity ensures reliability through its deterministic workflow—given the same inputs and parameters, it consistently generates identical results.
@@ 6 Conclusions
    By leveraging data from OSM, which is continuously updated by volunteers, VoxCity ensures its 3D city models remain relatively current. Many open 3D city models are not continuously updated and can quickly become obsolete; however, VoxCity mitigates this limitation by providing a more up-to-date representation of the urban form.
@@ Acknowledgements
We express our gratitude to the members of the NUS Urban Analytics Lab for the valuable discussions. This research has been supported by Takenaka Corporation. This research is part of the projects (i) Large-scale 3D Geospatial Data for Urban Analytics, which is supported by the National University of Singapore under the Start Up Grant; and (ii) Multi-scale Digital Twins for the Urban Environment: From Heartbeats to Cities, which is supported by the Singapore Ministry of Education Academic Research Fund Tier 1. We would like to thank the Singapore International Graduate Award (SINGA) scholarship provided by the Agency for Science, Technology, and Research (A*STAR), and NUS Research Scholarship and the President’s Graduate Fellowship (PGF) provided by NUS.
@@
The geospatial data used in this paper is openly available (see Section 3).
Code availability
The source code of VoxCity is openly available on GitHub (https://github.com/kunifujiwara/VoxCity), with comprehensive documentation (https://voxcity.readthedocs.io/en/latest/), a demo Google Colab notebook (https://colab.research.google.com/drive/1Lofd3RawKMr6QuUsamGaF48u2MN0hfrP?usp=sharing), and a tutorial video (https://youtu.be/qHusvKB07qk). VoxCity is distributed under the MIT License, a permissive open-source license that allows for free use, modification, and distribution, including for commercial purposes, with minimal restrictions. The package is available via PyPI (pip install voxcity).
Users can report issues, request features, and contribute code through GitHub. We welcome community contributions through pull requests. We plan to provide long-term support for stable releases, ensuring the package remains functional as underlying data sources and dependencies evolve.
Appendix A Environmental simulations and urban elements
Table 7: Assessment categories for environmental simulation and their corresponding urban elements
Assessment categories References Urban elements
Thermal comfort / heat stress Chen et al. [41], Yuan et al. [2], Lindberg et al. [1], Kong et al. [42], Li and Wang [43], Galal et al. [54], Zhang et al. [55], Hong and Lin [56], Wu et al. [57], Morakinyo and Lam [58], Li et al. [59], Wang et al. [60], Robitu et al. [66], Du et al. [67], Imam Syafii et al. [68] Buildings, trees, land cover, terrain elevation
Sunlight exposure / photovoltaic potential Grant et al. [53], Fath et al. [172] Buildings, trees
Wind comfort / wind energy potential / wind pressure Blocken et al. [49], Mou et al. [51], Hong and Lin [56] Buildings, trees
Building energy performance Forouzandeh [44], Gros et al. [6], Bouyer et al. [7] Buildings, land cover (low vegetation, etc.)
Air quality Kwak et al. [48] Buildings, terrain elevation, air pollutant emissions
Visual comfort / green view index / sky view factor Oh [5], Yi and Kim [72], Yu et al. [71], Labib et al. [74], Fujiwara et al. [73] Buildings, trees, land cover, terrain elevation
Appendix B Data structure of voxel city models
Figure 13: 3-dimensional array representing semantic voxel city models.
Appendix C 3D models for diverse geographies worldwide
Figure 14: Additional examples of output 3D city models from VoxCity. The 3D rendering was performed in VoxCity’s visualization module. BH, TH, LC, and TE represent building height, tree canopy height, land cover, and terrain elevation, respectively.
Appendix D Calculation of solar irradiance
We implemented a calculation combining the methods by Pružinec and Ďuračiová [143] and Monsi and Saeki [144]. Pružinec and Ďuračiová [143] introduced a ray-tracing and voxel-based method, while Monsi and Saeki [144] introduced a method to calculate transmittance of trees from leaf area density (LAD). For instantaneous values, the module first calculates solar azimuth and elevation angles using the location (longitude and latitude) of the target area and the target timestamp, employing a Python package ‘astral’ [^14] https://github.com/sffjunkie/astral. Direct and diffuse irradiance are then calculated using a ray-tracing technique, incorporating the computed sun position as well as direct normal and diffuse horizontal irradiance that is not affected by shading. In the ray-tracing process, a transmittance determined by the Beer-Lambert law (Equation 1) is incorporated when rays pass through tree voxels:
[…]
For cumulative values, the module calculates instantaneous irradiance at each timestamp from the start to the end of a specified period and then sums the results. To incorporate locally relevant climate conditions (including direct normal and diffuse horizontal irradiance), the module can use EnergyPlus Weather (EPW) files. Users may supply their own EPW files; otherwise, the module can automatically download the nearest EPW file from Climate.OneBuilding.Org  [^15] https://climate.onebuilding.org/.
Appendix E Time requirements for 3D city model generation and simulations
Table 8: Time requirements for 3D city model generation and simulations using VoxCity for an area in New York City. The workflow includes both manual user inputs (specifying parameters, target areas, etc.) and automated computational processes (data downloading, ray-tracing calculations, etc.). The target area, data sources, and simulation settings are the same as those in Figures 7 and 8. Processing was performed on our system with an Intel Core i9-13900 processor and an internet connection with a measured download speed of 180 Mbps.
Process Type Breakdown Time required
Model generation Manual Specifying target area, setting data sources and parameters 90 sec
Computational Downloading data sources, generating a voxel city model 40 sec
Solar irradiance simulation Manual Setting target time range 30 sec
Computational Downloading the nearest weather file, ray-tracing calculation (for 8760 time steps) 50 sec
Green view index calculation Manual - 0 sec
Computational Ray-tracing calculation 1 sec
Landmark visibility analysis Manual Specifying landmark buildings 30 sec
Computational Ray-tracing calculation 1 sec
Appendix F ENVI-met Material Mapping and Parameters
Table 9: Mapping of voxel classes in VoxCity to ENVI-met material and vegetation IDs with their basic properties. The model uses a layered structure where vegetation and building objects are placed on top of ground/soil profiles (e.g., rangeland = grass on sandy loam; wet land = grass on water; building = wall on sandy loam).
Voxel classes in
VoxCity
Material type ENVI-met material/
vegetation ID
Basic properties
Bareland, Rangeland, Shrub, Moss and lichen, Agriculture land, Tree, Snow and ice, Building, No Data Ground/soil 000000 (default sandy loam) Albedo: 0.20, Emissivity: 0.98, Roughness: 0.015 m, Sandy loam soil profile (19 layers)
Wet land, Mangroves, Water Ground/soil 0200WW (deep water) Albedo: 0.00, Emissivity: 0.96, Roughness: 0.01 m, Water mixing coeff: 0.001, Turbidity: 2.1
Road Ground/soil 0200ST (asphalt road) Albedo: 0.12, Emissivity: 0.90, Roughness: 0.01 m, 9 layers asphalt over 10 layers sandy loam
Rangeland, Moss and lichen, Agriculture land, Wet land Vegetation 0200XX (grass 25 cm aver. dense) Height: 0.25 m, Albedo: 0.20, Emissivity: 0.97, Transmittance: 0.30, rs_min: 200 s/m, LAD: uniform 0.30 m²/m³, Root depth: 0.20 m
Shrub Vegetation 0200H1 (hedge dense, 1 m) Height: 1.00 m, Albedo: 0.20, Emissivity: 0.97, Transmittance: 0.30, rs_min: 400 s/m, LAD: uniform 1.00 m²/m³, Root depth: 1.00 m
Tree Vegetation - (user-defined 3D plants by canopy height) Albedo: 0.18, Transmittance: 0.30, rs_min: 0.0 s/m, LAD: 1.0 m-1, Wood density: 690 kg/m³
Building Building wall 000000 (default wall – moderate insulation) 3-layer (plaster/insulation/concrete), Total thickness: 0.31 m, Roughness: 0.02 m, Can be greened